# Mask Approximation Net: A Novel Diffusion Model Approach for Remote Sensing Change Captioning

### v2

@@ Abstract
Remote sensing image change description represents an innovative multimodal task within the realm of remote sensing processing. This task not only facilitates the detection of alterations in surface conditions, but also provides comprehensive descriptions of these changes, thereby improving human interpretability and interactivity. Generally, existing deep-learning-based methods predominantly utilized a three-stage framework that successively perform feature extraction, feature fusion, and localization from bitemporal images before text generation. However, this reliance often leads to an excessive focus on the design of specific network architectures and restricts the feature distributions to the dataset at hand, which in turn results in limited generalizability and robustness during application. To address these limitations, this paper proposes a novel approach for remote sensing image change detection and description that incorporates diffusion models, aiming to transition the emphasis of modeling paradigms from conventional feature learning to data distribution learning. The proposed method primarily includes a simple multi-scale change detection module, whose output features are subsequently refined by an well-designed diffusion model. Furthermore, we introduce a frequency-guided complex filter module to boost the model performance by managing high-frequency noise throughout the diffusion process. We validate the effectiveness of our proposed method across several datasets for remote sensing change detection and description, showcasing its superior performance compared to existing techniques. The code will be available at MaskApproxNet after a possible publication.
@@ I Introduction
Remote sensing change detection is a vital task in analyzing and monitoring temporal variations on the Earth’s surface through satellite or aerial imagery. By comparing images acquired at different time intervals, this task aims to identify changes such as land use modifications [1], deforestation, urbanization [2], and environmental degradation [3; 4]. Traditionally, change detection [5; 6] has relied on pixel-level or object-level approaches to highlight areas with significant differences between bi-temporal images, providing essential information for environmental monitoring, disaster management, and urban planning.
@@
However, while these methods are effective at detecting the presence or absence of changes, they often lack the ability to describe the nature and significance of the detected changes in a human-readable format [7]. This limitation highlights the need for remote sensing change captioning.
@@
In recent years, remote sensing change captioning has garnered significant attention due to its capability to provide detailed, natural-language descriptions of temporal variations on the Earth’s surface [8]. Despite advances in this area, existing deep learning-based methods that are predominantly based on Convolutional Neural Networks (CNNs) exhibit several shortcomings, hindering their practical applicability [9; 10]. A primary concern lies in their reliance on complex network architectures specifically designed for feature extraction from bi-temporal images, as shown in Figure 1 (a). This often results in limited generalizability across datasets, as such methods struggle to adapt to different domains and their effectiveness diminishes when applied to scenarios or datasets that diverge significantly from those used during training. Moreover, CNN-based methods are prone to overfitting, making it difficult to ensure robustness and reliability in real-world applications, especially where data distribution varies. These issues significantly restrict the practical use of remote sensing change captioning in diverse and dynamic environments.
@@
To address the above-mentioned limitations, in this article, we propose a novel method by leveraging diffusion models for remote sensing image change captioning [11]. As illustrated in Figure 1, our approach diverges from previous methods that depend on complex attention networks to generate change masks. Instead, we utilize diffusion models, which prioritize learning data distributions over mere feature extraction. This paradigm shift enables the model to better capture the temporal complexities inherent in remote sensing imagery. Furthermore, we integrate a multi-scale change captioning module that identifies changes at various scales and refines the outputs through a diffusion process, resulting in more robust and adaptive caption generation. The proposed strategy not only simplifies the entire pipeline for change detection and captioning but also enhances generalizability and robustness by eliminating the need for intricate network designs. Furthermore, we introduce a frequency-guided complex filter module that effectively manages high-frequency noise, ensuring that the learned representations remain accurate and resilient, thereby improving the overall quality of the generated captions.
@@
  • Introduction of the Diffusion Model: We propose the first application of the diffusion model for remote sensing change captioning, enhancing performance and streamlining the task pipeline.
@@ II-A Change Captioning
The RSICC task [12] has garnered significant attention in recent years due to its capability to describe differences between bitemporal remote sensing (RS) images using natural language. Park et al. [13] presents a novel Dual Dynamic Attention Model (DUDA) to perform robust change captioning. Hoxha et al. [14] introduced early and late feature fusion strategies to integrate bitemporal visual features, utilizing an RNN and a multiclass SVM decoder for generating change captions. Chouaf et al. [15] were among the first to explore the RSICC task, employing a CNN as a visual encoder to capture temporal scene changes and an RNN as a decoder to produce descriptive change captions.
Transformer networks [16], incorporating the multi-head attention (MHA) mechanism, have gained prominence in image analysis and achieved notable success in image change captioning. Liu et al. [17] advanced the field by proposing a transformer-based encoder-decoder framework for RSICC. Their method includes a dual-branch transformer encoder for detecting scene changes and a multistage fusion module for combining multilayer features to generate change descriptions. Subsequently, Liu et al. [18] refined their approach by introducing progressive difference perception transformer layers, which effectively capture both high-level and low-level semantic changes. Furthermore, Liu et al. [19] proposed a prompt-based strategy that leverages pretrained large language models (LLMs) for RSICC tasks, using visual features, change classes, and language representations as prompts for a frozen LLM to generate change captions. Chang et al. [20] developed an attentive network for RS change captioning, named Chg2Cap, which harnesses the strengths of transformer models commonly used in NLP. Sun et al. [21] proposes a lightweight Sparse Focus Transformer for remote sensing image change captioning, which significantly reduces parameters and computational complexity while maintaining competitive performance. In DiffusionRSCC [22] introduced an innovative diffusion model for RSICC, employing a forward noising and reverse denoising process to learn the probabilistic distribution of the input. Zhou et al.[23] proposed a bitemporal pretraining method that leverages self-supervised learning on a large-scale bitemporal RS image dataset. This approach reduces data distribution and input gaps, resulting in more suitable features for RSICC and improved model generalization.
@@ II-B Generative Models
Within the enormous models for CD, DDPM-based [27] architectures emerge with distinguished advantages over traditional CNNs and transformers. In DDPM-CD [28] by pre-training the DDPM on a large set of unlabeled remote sensing images, multi-scale feature representations are obtained. A lightweight change detection classifier is then fine-tuned to detect precise changes. Wen et al. [29] designed to guide the generation of change detection maps by exploiting multi-level difference features. The Similar ideas have also applied in medicine field. Yu et al. [22] employs a noise predictor conditioned on cross-modal features to generate human-like descriptions of semantic changes between bi-temporal remote sensing image pairs. Wu et al. [30] proposes dynamic conditional encoding to enhance step-wise regional attention to mitigate high-frequency noise for medical image segmentation. Zhang et al. [31] presents a graph attention-guided diffusion model tailored for liver vessel segmentation. By integrating graph attention mechanisms with diffusion probabilistic models, the approach effectively captures complex vascular structures in liver images.
@@ III Methodology
Figure 2: The overall framework of the proposed method, which consists of two main stages, i.e., the mask approximation phase for feature processing and change feature generation, and the text decoding phase for producing change detection descriptions.
@@ III-B Mask Approximation Phase
Existing methods typically rely on multi-stage architectures, where each stage handles a specific task, such as feature extraction, feature fusion or location, and description generation. While these methods have shown some success, they often suffer from inefficiencies and complexity, as the intermediate outputs from each stage may not always align well, leading to the propagation of errors and unnecessary computational overhead.
@@ III-B1 Mask Approx Module
$X_{i}^{\text{pre}}=\text{ResNet}(I_{\text{pre}},i),\quad X_{i}^{\text{post}}=\text{ResNet}(I_{\text{post}},i).$
@@
$\overline{X}=\text{Conv}(\text{DifferenceModule}(X_{i}^{\text{pre}},X_{i}^{\text{post}})).$
@@
$\overline{X}_{\text{upsampled}}=\text{TransposeConv}(\overline{X}),$
@@
$\overline{X}_{\text{refined}}=\text{ResidualBlock}(\overline{X}_{\text{upsampled}}).$
@@
$\text{DiffImage}=\text{Conv}(\overline{X}_{\text{refined}}).$
@@
$\bm{\epsilon_{\theta}}(\mathbf{x}_{t},I_{\text{pre}},I_{\text{post}},t)=\text{UNet}($ $\text{Cat}(I_{\text{pre}},\text{MaskNoise}),$ (7)
$\text{Cat}(I_{\text{post}},\text{MaskNoise}),$
@@ III-B2 Frequency Noise Filter
$W_{c}(f_{x},f_{y})=\exp\left(-\alpha\cdot\|F(f_{x},f_{y})\|_{2}^{2}\right),$
@@
$A_{c}=\sigma\left(W_{a}\cdot\text{GAP}(X_{c}^{\prime})+b_{a}\right),$
@@
$X_{FGCF}=\sum_{c=1}^{C}A_{c}\cdot X_{c}^{\prime}.$
@@
III-C Text Decoding Phase
$\bm{T}_{\text{embed}}=E_{\text{embed}}(\bm{t})+E_{\text{pos}},$ (9)
where $E_{\text{embed}}(\bm{t})$ represents the token embeddings and $E_{\text{pos}}$ is the positional encoding, which captures the sequential order of the tokens. These embeddings serve as the initial input to the transformer decoder as shown in Figure 5.
Figure 5: The diagram illustrates the algorithmic workflow of the transformer decoder, highlighting the key processes of embedding, masked multi-head attention, feed-forward computation, and autoregressive token generation.
The next step involves passing $\bm{T}_{\text{embed}}$ through the masked multi-head self-attention (MHA) mechanism. At the core of this operation lies the computation of attention heads, defined as:
where the attention function is expressed as,
$=$ $\mathrm{Concat}(\bm{Head}_{1},\dots,\bm{Head}_{h})\cdot\bm{W}^{O}.$
Following the MHA sub-layer, the output undergoes further processing through a feed-forward network (FFN). The final output of the decoder layer is obtained by incorporating the input embedding $\bm{T}_{\text{embed}}^{i-1}$ with the output of the FFN via residual connections:
IV Experimental
IV-A Datasets
@@
IV-B Experimental Setup
IV-B1 Evaluation Metrics
  • BLEU (Bilingual Evaluation Understudy): Introduced by Papineni et al. [36], BLEU is a precision-based metric for evaluating machine translation quality. It calculates the overlap between n-grams of the candidate and reference translations, with n typically ranging from 1 to 4. The BLEU score is computed as:
    $\text{BLEU}=\text{BP}\times\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right),$ (15)
    where $p_{n}$ is the precision of n-grams, $w_{n}$ are positive weights summing to one (often uniform), and BP is the brevity penalty to penalize short translations:
    $\text{BP}=\begin{cases}1,&\text{if }c>r,\\ \exp(1-\frac{r}{c}),&\text{if }c\leq r.\end{cases}$ (16)
    Here, $c$ is the length of the candidate translation, and $r$ is the length of the reference translation.
  • ROUGE-L (Recall-Oriented Understudy for Gisting Evaluation): Part of the ROUGE metric family introduced by Lin [37], ROUGE-L evaluates the quality of summaries by measuring the longest common subsequence (LCS) between the candidate and reference texts. The ROUGE-L score considers both precision and recall:
    $F_{\beta}=\frac{(1+\beta^{2})\times\text{Precision}\times\text{Recall}}{\text{Precision}+\beta^{2}\times\text{Recall}},$ (17)
    where Precision = $\frac{\text{LCS}(X,Y)}{\text{length of }X}$, Recall = $\frac{\text{LCS}(X,Y)}{\text{length of }Y}$, and $\beta$ is a parameter that determines the relative importance of precision and recall.
  • METEOR (Metric for Evaluation of Translation with Explicit ORdering): Proposed by Banerjee and Lavie [38], METEOR evaluates translation quality by aligning words in the candidate and reference translations using exact matches, stemming, and synonyms. It computes a harmonic mean of unigram precision and recall, giving higher weight to recall, and includes a penalty for fragmented matches:
    $\text{METEOR}=F_{mean}\times(1-\text{Penalty}),$ (18)
    where $F_{mean}=\frac{10\times\text{Precision}\times\text{Recall}}{\text{Recall}+9\times\text{Precision}}$ and Penalty = $0.5\times\left(\frac{\text{chunks}}{\text{matches}}\right)^{3}$.
  • CIDEr-D (Consensus-based Image Description Evaluation): Introduced by Vedantam et al. [39], CIDEr-D is designed for evaluating image descriptions by measuring the consensus between a candidate description and a set of reference descriptions. CIDEr-D applies TF-IDF weighting to n-grams, emphasizing the importance of less frequent, informative n-grams:
[…]
    where $m$ is the number of reference sentences, $w_{n}$ is the TF-IDF weight for n-grams, and $\text{CIDEr}_{n}$ measures the similarity of n-grams between the candidate and reference descriptions.
IV-B2 Experimental Details
IV-C Quantitative Results
@@
Figure 6: Visualized image and captioning examples generated by Mask Apporx Net on the LEVIR-MCI dataset.
@@ IV-D Qualitative Visualization
Figure 7: Visualized image and captioning examples generated by Mask Apporx Net on the WHU-CDC dataset
@@
ID Frequency Noise Filter BLEU-1 BLEU-2 BLEU-3 BLEU-4
0 85.63 76.79 70.41 63.96
1 85.90 77.12 70.72 64.32

### v3

@@ Abstract
Remote sensing image change description represents an innovative multimodal task within the realm of remote sensing processing. This task not only facilitates the detection of alterations in surface conditions, but also provides comprehensive descriptions of these changes, thereby improving human interpretability and interactivity. Current deep learning methods typically adopt a three-stage framework consisting of feature extraction, feature fusion, and change localization, followed by text generation. Most approaches focus heavily on designing complex network modules but lack solid theoretical guidance, relying instead on extensive empirical experimentation and iterative tuning of network components. This experience-driven design paradigm may lead to overfitting and design bottlenecks, thereby limiting the model’s generalizability and adaptability. To address these limitations, this paper proposes a paradigm that shift towards data distribution learning using diffusion models, reinforced by frequency-domain noise filtering, to provide a theoretically motivated and practically effective solution to multimodal remote sensing change description. The proposed method primarily includes a simple multi-scale change detection module, whose output features are subsequently refined by a well-designed diffusion model. Furthermore, we introduce a frequency-guided complex filter module to boost the model performance by managing high-frequency noise throughout the diffusion process. We validate the effectiveness of our proposed method across several datasets for remote sensing change detection and description, showcasing its superior performance compared to existing techniques. The code will be available at MaskApproxNet.
@@ I Introduction
Remote sensing change detection is a vital task in analyzing and monitoring temporal variations on the Earth’s surface through satellite or aerial imagery. By comparing images acquired at different time intervals, this task aims to identify changes such as land use modifications [1], deforestation, urbanization [2], and environmental degradation [3, 4]. Traditionally, change detection [5, 6, 7, 8] has relied on pixel-level or object-level approaches to highlight areas with significant differences between bi-temporal images [9, 10], providing essential information for environmental monitoring, disaster management[11, 12, 13], and urban planning. However, while these methods are effective at detecting the presence or absence of changes, they often lack the ability to describe the nature and significance of the detected changes in a human-readable format [14, 15]. This limitation highlights the need for remote sensing change captioning.
@@
In recent years, remote sensing change captioning has garnered significant attention due to its capability to provide detailed, natural-language descriptions of temporal variations on the Earth’s surface [16, 17]. Despite advances in this area, existing deep learning-based methods [18, 19, 20] that are predominantly based on Convolutional Neural Networks (CNNs) exhibit several shortcomings, hindering their practical applicability [21, 22]. A primary concern lies in their reliance on a three-stage framework—feature extraction, feature fusion, and change localization—followed by text generation, where the design of complex network modules is mainly driven by empirical experimentation rather than solid theoretical foundations, as illustrated in Figure 1 (a). This experience-driven approach often leads to overfitting and design bottlenecks, restricting the models’ robustness and adaptability. Consequently, their generalizability across diverse datasets and application scenarios remains limited, especially when faced with domains that differ significantly from the training data.However, diffusion models offer a promising alternative. As a likelihood-based generative approach, diffusion models employ a Markov chain to progressively add noise to data and learn to reconstruct samples by reversing this noising process. This stochastic process inherently captures data diversity and complexity, which effectively reduces overfitting [23, 24] and enhances model generalization. Unlike conventional CNN-based methods that are prone to overfitting and often struggle with robustness and reliability in real-world, dynamic remote sensing scenarios, diffusion models are better suited to handle the diverse and complex data distributions encountered in change captioning tasks, thereby improving practical applicability and performance.
@@
To address the above-mentioned limitations, in this article, we propose a novel method by leveraging diffusion models for remote sensing image change captioning [25]. As illustrated in Figure 1 (b), our approach diverges from previous methods that depend on complex attention networks to generate change masks. Instead, we utilize diffusion models, which prioritize learning data distributions over mere feature extraction. Diffusion models have shown remarkable success in both continuous visual data [26, 27] and discrete language data domains [28, 29], producing high-quality samples while deeply capturing connections between pixels, words, and image-text relationships in multi-modal tasks [30, 31]. By incorporating a paradigm shift that enhances their ability to grasp temporal complexities, these models are particularly adept at handling the dynamic intricacies of remote sensing imagery. Furthermore, we integrate a multi-scale change captioning module that identifies changes at various scales and refines the outputs through a diffusion process, resulting in more robust and adaptive caption generation. The proposed strategy not only simplifies the entire pipeline for change detection and captioning but also enhances generalizability and robustness by eliminating the need for intricate network designs. Besides, we introduce a frequency-guided complex filter module that effectively manages high-frequency noise, ensuring that the learned representations remain accurate and resilient, thereby improving the overall quality of the generated captions.
@@
  1. 1.
    Introduction of the Diffusion Model: To the best of our knowledge, our work is the first to introduce diffusion models specifically into the image processing component of the remote sensing image change captioning task. By leveraging the strong theoretical foundation of diffusion models to effectively model complex temporal variations in multi-temporal remote sensing imagery, our approach aims to improve performance while simplifying the overall task framework.
  2. 2.
@@ II-A Change Captioning
The Remote Sensing Image Change Captioning(RSICC) task [32] has garnered significant attention in recent years due to its capability to describe differences between bitemporal remote sensing (RS) images using natural language. Park et al. [33] presents a novel Dual Dynamic Attention Model (DUDA) to perform robust change captioning. Hoxha et al. [34] introduced early and late feature fusion strategies to integrate bitemporal visual features, utilizing an RNN and a multiclass SVM decoder for generating change captions. Chouaf et al. [35] were among the first to explore the RSICC task, employing a CNN as a visual encoder to capture temporal scene changes and an RNN as a decoder to produce descriptive change captions.
Figure 2: The overall framework of the proposed method includes two main stages. The first stage employs a conditional diffusion model that generates change detection (CD) maps by progressively denoising Gaussian noise, guided by the pre-change and post-change images ($I_{pre}$ and $I_{post}$). This iterative process allows effective modeling of complex temporal changes. The second stage decodes the generated CD maps into descriptive text, completing the remote sensing change captioning task.
Transformer networks [36], incorporating the multi-head attention (MHA) mechanism, have gained prominence in image analysis and achieved notable success in image change captioning. Liu et al. have made a series of significant contributions to RSICC. They first proposed a transformer-based encoder-decoder framework featuring a dual-branch transformer encoder for scene change detection and a multistage fusion module to integrate multilayer features for generating change descriptions [37]. Building on this, they introduced progressive difference perception transformer layers to better capture both high-level and low-level semantic changes [38]. More recently, they developed a prompt-based approach that leverages pretrained large language models (LLMs), using visual features, change classes, and language representations as prompts to guide a frozen LLM in generating change captions [39]. Kuckreja et al. [40] expanded the image-text pairs of existing different remote sensing datasets, a multifunctional model Geo Chat was proposed to generate change captions. Chang et al. [41] developed an attentive network for RS change captioning, named Chg2Cap, which harnesses the strengths of transformer models commonly used in NLP. Sun et al. [42] proposes a lightweight Sparse Focus Transformer for remote sensing image change captioning, which significantly reduces parameters and computational complexity while maintaining competitive performance. In DiffusionRSCC [43] introduced an innovative diffusion model for RSICC, employing a forward noising and reverse denoising process to learn the probabilistic distribution of the input. Zhou et al.[44] proposed a bitemporal pretraining method that leverages self-supervised learning on a large-scale bitemporal RS image dataset. This approach reduces data distribution and input gaps, resulting in more suitable features for RSICC and improved model generalization.
@@ II-B Generative Models
Within the enormous models for CD, Denoising Diffusion Probabilistic Model(DDPM) based [48] architectures emerge with distinguished advantages over traditional CNNs and transformers. In DDPM-CD [49] by pre-training the DDPM on a large set of unlabeled remote sensing images, multi-scale feature representations are obtained. A lightweight change detection classifier is then fine-tuned to detect precise changes. Wen et al. [50] designed to guide the generation of change detection maps by exploiting multi-level difference features. The Similar ideas have also applied in medicine field. Yu et al. [43] employs a noise predictor conditioned on cross-modal features to generate human-like descriptions of semantic changes between bi-temporal remote sensing image pairs. Zhang et al. [51] presents a graph attention-guided diffusion model tailored for liver vessel segmentation. By integrating graph attention mechanisms with diffusion probabilistic models, the approach effectively captures complex vascular structures in liver images.
@@ III-B Mask Approximation Phase
Existing methods usually rely on multi-stage architectures handling feature extraction, fusion, localization, and description separately. This leads to inefficiencies, as numerous stages cause misalignment between image features and text, propagating errors. Moreover, the complexity arises from empirically designed network modules lacking strong theoretical support, limiting model robustness and scalability.
@@ III-B1 Mask Approx Module
$X_{i}^{\text{pre}}=\text{ResNet}(I_{\text{pre}},i),\quad X_{i}^{\text{post}}=\text{ResNet}(I_{\text{post}},i).$ (6)
@@
$\overline{X}=\text{Conv}(\text{DifferenceModule}(X_{i}^{\text{pre}},X_{i}^{\text{post}})).$ (7)
@@
$\overline{X}_{\text{upsampled}}=\text{TransposeConv}(\overline{X}),$ (8)
@@
$\overline{X}_{\text{refined}}=\text{ResidualBlock}(\overline{X}_{\text{upsampled}}).$ (9)
@@
$\text{DiffImage}=\text{Conv}(\overline{X}_{\text{refined}}).$ (10)
@@
$\bm{\epsilon_{\theta}}(\mathbf{x}_{t},I_{\text{pre}},I_{\text{post}},t)=\text{UNet}($ $I_{\text{pre}},I_{\text{post}},$ (12)
@@ III-B2 Frequency Noise Filter
$W_{c}(f_{x},f_{y})=\exp\left(-\alpha\cdot\|F(f_{x},f_{y})\|_{2}^{2}\right),$ (15)
@@
$A_{c}=\sigma\left(W_{a}\cdot\text{GAP}(X_{c}^{\prime})+b_{a}\right),$ (18)
@@
$X_{FGCF}=\sum_{c=1}^{C}A_{c}\cdot X_{c}^{\prime}.$ (19)
@@
III-C Text Decoding Phase
$\bm{T}_{\text{embed}}=E_{\text{embed}}(\bm{t})+E_{\text{pos}},$ (20)
where $E_{\text{embed}}(\bm{t})$ represents the token embeddings and $E_{\text{pos}}$ is the positional encoding, which captures the sequential order of the tokens. These embeddings serve as the initial input to the transformer decoder.
Next, $\bm{T}_{\text{embed}}$ is processed via the masked multi-head self-attention (MHA) mechanism, where attention heads are computed as:
where the attention function is expressed as,
$=$ $\mathrm{Concat}(\bm{Head}_{1},\dots,\bm{Head}_{h})\cdot\bm{W}^{O}.$
Following the MHA layer, the output undergoes further processing through a feed-forward network (FFN). The final output of the decoder layer is obtained by incorporating the input embedding $\bm{T}_{\text{embed}}^{i-1}$ with the output of the FFN via residual connections:
IV Experimental
IV-A Datasets
@@
IV-B Experimental Setup
IV-B1 Evaluation Metrics
  • BLEU: A precision-based metric for machine translation that measures n-gram overlap between candidate and reference texts. A brevity penalty is used to penalize short outputs [56].
  • ROUGE-L: Evaluates summary quality based on the longest common subsequence (LCS) between candidate and reference, considering both precision and recall [57].
  • METEOR: Aligns words using exact matches, stemming, and synonyms; computes a harmonic mean of precision and recall with a fragmentation penalty [58].
  • CIDEr-D: Designed for image captioning, this metric compares n-gram similarity between candidate and reference descriptions using TF-IDF weighting [59].
IV-B2 Experimental Details
IV-C Quantitative Results
@@ IV-D Qualitative Visualization
Figure 5: Visualized image and captioning examples generated by Mask Apporx Net on the LEVIR-MCI (THE LEFT) and WHU-CDC (THE RIGHT) dataset
@@
FGCF BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr-D
- 85.63 76.79 70.41 63.96 39.70 75.53 137.24
85.90 77.12 70.72 64.32 39.91 75.67 137.71
@@
TABLE IV: Comparison of Model Parameters and Computational Cost (FLOPs) Across Different Methods with $256\times 256$ image size
ID Method Paramters(M) FLOPs
1 Capt-Rep-Diff 45.6 19.8 G
2 Capt-Att 46.1 19.9 G
3 Capt-Dual-Att 48.1 20.0 G
4 DUDA 52.8 20.3 G
5 MCCFormer-S 135.0 25.1 G
6 MCCFormer-D 135.0 25.1 G
7 RSICCFormer 145.3 27.1 G
8 PSNet 231.5 13.8 G
9 Prompt-CC 196.3 19.9 G
10 Chg2Cap 286.5 232.1 G
11 Sparse Focus 117.5 21.5 G
12 SEN 39.9 10.9 G
13 DiffusionRSCC 103.0 15.2 T
14 Mask Apporx Net(ours) 202.5 12.3 T